# Influence-aware Causal Autoencoder Network for Node Importance Ranking in Complex Networks

### v1

@@ II-B Assumptions
(Causal graph [25, 26]) For Bayesian Network $U,\mathbb{G},P>$, where $U$ denotes the set of variables, $\mathbb{G}$ is a Directed Acyclic Graph (DAG) on $U$ and $P$ denotes the probability distribution of $U$, then Bayesian Network $U,\mathbb{G},P>$ can be used to express the causal relationships between the variables in $U$. In DAG $\mathbb{G}$, for a pair of directly connected parent-child variables, the parent variable is the direct cause of the child variable, and the child variable is the direct outcome of the parent variable. We assume that the causality between variables can be expressed by $\mathbb{G}$ as a causal graph.
@@
(Causal Markov [27, 28]) A variable $X$ is independent of every other variable (except X’s effects) conditional on all of its direct causes.
@@
(Faithfulness [25, 29]) Every conditional independence that holds in the distribution $P$ is entailed by the Markov condition applied to $\mathbb{G}$. Faithfulness ensures that all conditional independencies in the data correspond to missing edges in the causal graph.
Figure 1: The framework of ICAN.
III Proposed method
In this section, we first show the overall framework of the proposed method, the Influence-aware Causal Autoencoder Network for node importance ranking (ICAN), and then describe its components in detail.
III-A Overview of the method
The framework of ICAN is illustrated in Fig. 1. ICAN consists of two core modules: the causal representation learning module and the causal ranking prediction module. The causal representation learning module aims to learn low-dimensional node embeddings that are causally related to node importance. Under Assumption 1, the causality among node embeddings can be described by a DAG $\mathbb{G}$. Upon obtaining the $p$-dimensional $(p\leq d)$ feature representation, the node influence score variable $\bm{Y}$, which can be used to characterize the node importance, is introduced as an additional node to form a new hidden layer. The adjacency matrix $\bm{W}=\{w_{ij}\}_{(p+1)\times(p+1)}$ associated with $\mathbb{G}$, which encodes the causal relationships among the low-dimensional node embeddings $\bm{H}^{(l)}$, is then integrated into the autoencoder framework to enable message passing under causal mechanisms. By optimizing $\bm{W}$, ICAN can learn optimal representations that are causally linked to node importance. The learned latent representations are then fed into the causal ranking prediction module, where the Markov Blanket (MB) of the node importance score variable is leveraged to predict the ranking. In this way, ICAN can produce robust, network-invariant node embeddings, enabling cross-network importance ranking predictions. In the following, we provide a detailed description of the ICAN framework.
III-B Causal representation learning module
The goal of the causal representation learning module is to infer causal relationships among low-dimensional node embeddings. In general, this module consists of two main components: an encoder and a decoder. We first generate a $d$-dimensional feature matrix $\bm{X}$ from the graph adjacency matrix $\bm{A}$ using Node2vec. The encoder, typically a Graph Convolutional Network (GCN), then takes both $\bm{X}$ and $\bm{A}$ as input. It learns to map this input into a $p$-dimensional low-dimensional representation. The decoder then uses this learned representation to reconstruct the original adjacency matrix. This integrated process of encoding and decoding can be summarized as follows:
@@ III-B Causal representation learning module
Influence-aware causal representation learning mechanism. Based on Assumption 1, learning the causal relationships among variables can be formulated as learning the matrix $\bm{W}$ associated with the DAG $\mathbb{G}$. To obtain node embeddings that are causally related to node importance, we introduce a pseudo-label variable $\bm{Y}$, termed the node influence score, which characterizes node importance based on the SIR model—a classical epidemic framework describing a spreading process where each node can be in one of three states: susceptible (S), infected (I), or recovered (R). At each time step, an infected node infects its susceptible neighbors with probability $\gamma$ and recovers with probability $\delta$, after which it cannot be infected again. The process continues until no infected nodes remain. This model captures the key dynamics of contagion and recovery, which are also relevant to information diffusion and network resilience. Therefore, it provides a reliable basis for evaluating the spreading ability of nodes and for training deep models to understand real-world network dynamics. In our experiments, each node $v$ is initialized as the only infected node, while all others are susceptible. The final fraction of infected and recovered nodes when the process stabilizes is denoted as $F_{v}$, representing the spreading influence of node $v$. The average value of $F_{v}$ over 100 simulations is used as the influence score, denoted by $\bm{Y}_{v}$. Then we run the SIR model for each node in the network to obtain the influence score variable $\bm{Y}$.
@@
where the adjacent matrix $\bm{W}$ is required to satisfy the constraint in Eq. (18). $g_{1}$ and $g_{2}$ can be regarded as two nonlinear functions used to describe the encoding and decoding process respectively. Therefore, it is noted here that the causal representation learning module can be seen as a graph autoencoder.
@@
III-C Causal ranking prediction module
@@
Task-aware ranking loss function. We introduce a new ranking loss function, CausalListMLE, which can be viewed as an extension of ListMLE [30] for optimizing the overall ranking prediction module. This list-wise ranking loss directly maximizes the likelihood of the entire ranked permutation, rather than focusing on individual scores as in point-wise losses like MSE. By modeling the ranking distribution, CausalListMLE focuses on preserving the correct relative order among all items, and typically achieves better performance in ranking-based learning tasks. The causal ranking loss $\mathcal{L}_{2}$ is defined as
@@ III-D The objective function
To ensure that the learned $\bm{W}$ forms a DAG, this objective must satisfy the acyclicity constraint [31]:
@@
Conversely, if $\mathbb{G}$ is acyclic, Lemma 1 implies $q_{ii}^{(k)}=0$ for all $i$ and $k$. Thus, $\mathrm{tr}(\bm{Q}^{k})=0$, and therefore
@@
To solve above equality-constrained problem, the augmented Lagrangian method [32, 33] is employed to convert it into an unconstrained problem. After introducing the Lagrange multiplier and the penalty term, the objective function Eq. (III-D) is reformulated as the following augmented Lagrangian function:
@@
IV Experiment
In this section, experiments are performed on the real-world networks to validate the effectiveness of the proposed ICAN method in the node importance ranking problem. In the following, the experimental settings such as the used datasets, evaluation criterion, comparison methods and model parameters for the experiments are described first, followed by a detailed analysis of the results. To systematically evaluate ICAN, we define five research questions (RQs)—our “Five CANs”:
  • RQ1: Can ICAN learn features that are causally relevant to node importance from synthetic networks—without direct access to the target network structure—and generalize well across diverse target networks?
  • RQ2: Can the feature-task co-optimization mechanism in ICAN lead to improvement in the model’s performance?
  • RQ3: Can the causal ranking loss defined in ICAN contribute to the feature representation learning process?
  • RQ4: Can the model identify important nodes that are not simply high-degree hubs, capturing more nuanced influence patterns across the network?
  • RQ5: Can the model achieve consistently high performance on one give target network when trained on different types of generative networks?
IV-A Description of datasets
@@
Training datasets. Five representative synthetic network models are used for training, including Barabási–Albert (BA) [34, 35], Extreme Homogeneous (EH) [36], Erdös-Rényi (ER) random-graph [37], Q-Snapback (QS) [38] and Random Hexagon (RH) [38]. The BA networks are generated using the preferential attachment mechanism, where new nodes are sequentially added and connected to existing nodes with a probability proportional to their current degrees, leading to a scale-free degree distribution. The EH networks are obtained by performing additional random edge rectifications on ER networks to transform their degree distribution from Poisson to near-uniform. The ER random graphs are generated by connecting each pair of nodes independently with a fixed probability $p$, producing networks with a binomial degree distribution. The QS networks consist of a directed backbone chain with multiple probabilistic “snapback” edges linking newly added nodes to previously added ones within a defined range, thereby enhancing backward connectivity and robustness. The RH networks are composed of randomly connected hexagonal substructures that capture local clustering and spatial organization similar to lattice-like systems. All trained networks comprise 1,000 nodes with an average degree of 4.
@@
  • Karate [39]: A human social network, in which each node represents a member in the club and each link shows the relationship between two members.
  • Jazz [40]: A human collaboration network, in which each node represents a jazz musician and each link indicates that two musicians have performed together in the same band.
  • Email-univ [41]:An email communication network from the University Rovira i Virgili in Spain, where nodes represent users and edges indicate at least one email was exchanged.
  • USAir [42]: A directed weighted network represents the flight connections among 1,574 U.S. airports in 2010. Each node corresponds to an airport, and each directed edge indicates the existence of at least one flight from the origin to the destination airport. The edge weight quantifies the total number of flights operated on that route during the year.
  • Vidal [43]: A network represents an initial version of a proteomescale map of Human binary protein–protein interactions.
  • Email-dnc [44]: A directed unweighted network derived from the 2016 Democratic National Committee (DNC) email leak. Each node represents an individual user, and each directed edge indicates that at least one email was sent from one user to another.
The key statistical properties of the real-world networks are summed up in Table I . $n$ represents the number of nodes, $m$ is the number of edges, $\langle k\rangle$ represents the average node degree, $k_{\text{max}}$ is the maximum degree, $c$ stands for the average clustering coefficient, $AS$ indicates the degree assortativity, and $HE$ is the degree heterogeneity.
TABLE I: The statistical properties of the real-world networks.
[…]
Karate 34 78 4.5882 17 0 -0.1352 3.1064
Jazz 198 2742 27.7 100 0.5376 0.0474 3.1010
Email-univ 1133 5451 9.6 71 0.2201 0.0782 1.9421
USAir 1574 17215 21.9 314 0.5042 -0.1132 5.1303
Vidal 3023 6149 4.1 129 0.0658 -0.1256 3.7960
Email-dnc 2029 4465 4.4 404 0.1948 -0.3014 16.6476
Cora 2708 5263 3.8 168 0.2389 -0.0659 2.8035
IV-B Evaluation criterion
In this section, we introduce the evaluation metric used to measure the effectiveness of the predicted ranking results by ICAN, Kendall’s $\tau$ coefficient [45]. This non-parametric statistic is widely employed to measure the ordinal association between two ranking lists. It quantifies the similarity between two ranking orders by comparing the numbers of concordant and discordant pairs. The definition of $\tau$ is
$\tau=\frac{2(N_{c}-N_{d})}{n(n-1)}$},
IV-C Comparative methods
To demonstrate the advantages of the ICAN method in node ranking, we implement several state-of-the-art baselines. A brief summary of these methods is provided below.
  • Degree Centrality (DC) [12]. It quantifies the influence of a node based on the number of direct connections it has to other nodes in the graph. The degree centrality of a node u is formally defined as
@@ IV-C Comparative methods
  • Betweenness Centrality (BC) [46]. It measures the importance of a node by quantifying how frequently it appears on the shortest paths between all pairs of nodes in the network.
@@
  • Eigenvector Centrality (EC) [47]. It assigns each node a centrality score based not only on its own connectivity but also on the importance of the nodes it is connected to. The EC score $\bm{X}_{u}$ of node $u$ satisfies
@@
  • H-index (HI) [48]. The H-index of a node considers the degrees of its neighbors and reflects how many of them are themselves influential. For a node $u$ with neighbors $j_{1},j_{2},\ldots,j_{k}$, the H-index is defined as
@@
  • $K$-Shell (KS) [49]. The $K$-shell algorithm operates through an iterative pruning process. It iteratively peels away network layers to assign each node a $k$-value indicating its coreness based on residual connections. A higher $K$-shell value indicates greater node influence.
  • GNN-Bet [50]. This method enables node ranking by employing a dual-path aggregation mechanism that separately propagates features along incoming and outgoing shortest paths using row-wise modified adjacency matrices, with final ranking scores generated through multiplicative combination of the aggregated path representations.
  • GNN-Close [50]. This approach achieves node ranking through a hierarchical feature aggregation scheme that combines normal adjacency operations in the initial layer with column-wise modified matrices in subsequent layers to constrain feature propagation along shortest paths, producing final ranking scores via summation of layer-wise outputs.
  • RCNN [15]. This method leverages convolutional neural networks to extract features from node-centric subgraph structures and ranks node importance based on learned representations.
  • CGNN [19]. This method enhances RCNN by incorporating additional GNN layers, improving the capture of topological dependencies for more accurate vital node identification.
  • AGNN [16]. This method combines a GCN-based autoencoder with a GNN ranking head to jointly learn structural embeddings and predict node importance using listwise ranking loss.
IV-D Implementation details

### v2

@@ 3.2 Assumptions
(Causal graph [pearl2014probabilistic]) For Bayesian Network $U,\mathbb{G},P>$, where $U$ denotes the set of variables, $\mathbb{G}$ is a Directed Acyclic Graph (DAG) on $U$ and $P$ denotes the probability distribution of $U$, then Bayesian Network $U,\mathbb{G},P>$ can be used to express the causal relationships between the variables in $U$. In DAG $\mathbb{G}$, for a pair of directly connected parent-child variables, the parent variable is the direct cause of the child variable, and the child variable is the direct outcome of the parent variable. We assume that the causality between variables can be expressed by $\mathbb{G}$ as a causal graph.
@@
(Causal Markov [fields1990independence]) A variable $X$ is independent of every other variable (except X’s effects) conditional on all of its direct causes.
@@
(Faithfulness [guo2023adaptive]) Every conditional independence that holds in the distribution $P$ is entailed by the Markov condition applied to $\mathbb{G}$. Faithfulness ensures that all conditional independencies in the data correspond to missing edges in the causal graph.
4 Proposed method
In this section, we first show the overall framework of the proposed method, the Influence-aware Causal Autoencoder model for Node importance ranking in complex networks (ICAN), and then describe its components in detail.
4.1 Overview of the method
The framework of ICAN is illustrated in Fig. 1. ICAN consists of two core modules: the causal node embedding module and the causal ranking prediction module. The causal representation learning module aims to learn low-dimensional node embeddings that are causally related to node importance. Under Assumption 1, the causality among node embeddings can be described by a DAG $\mathbb{G}$. Upon obtaining the $p$-dimensional $(p\leq d)$ feature representation, the node influence score variable $\bm{Y}$, which can be used to characterize the node importance, is introduced as an additional node to form a new hidden layer. The adjacency matrix $\bm{W}=\{w_{ij}\}_{(p+1)\times(p+1)}$ associated with $\mathbb{G}$, which encodes the causal relationships among the low-dimensional node embeddings $\bm{H}^{(l)}$, is then integrated into the autoencoder framework to enable message passing under causal mechanisms. By optimizing $\bm{W}$, ICAN can learn optimal representations that are causally linked to node importance. The learned latent representations are then fed into the causal ranking prediction module, where the Markov Blanket (MB) of the node importance score variable is leveraged to predict the ranking. In this way, ICAN can produce robust, network-invariant node embeddings, enabling cross-network importance ranking predictions. In the following, we provide a detailed description of the ICAN framework.
Figure 1: The framework outline of ICAN.
4.2 Causal node embedding module
The goal of the causal node embedding module is to infer causal relationships among low-dimensional node representations. In general, this module consists of two main components: an encoder and a decoder. We first generate a $d$-dimensional feature matrix $\bm{X}$ from the graph adjacency matrix $\bm{A}$ using Node2vec. The encoder, typically a Graph Convolutional Network (GCN), then takes both $\bm{X}$ and $\bm{A}$ as input. It learns to map this input into a $p$-dimensional low-dimensional representation. The decoder then uses this learned representation to reconstruct the original adjacency matrix. This integrated process of encoding and decoding can be summarized as follows:
@@ 4.2 Causal node embedding module
Influence-aware causal node embedding mechanism. Based on Assumption 1, learning the causal relationships among variables can be formulated as learning the matrix $\bm{W}$ associated with the DAG $\mathbb{G}$. To obtain node embeddings that are causally related to node importance, we introduce a pseudo-label variable $\bm{Y}$, termed the node influence score, which characterizes node importance based on the SIR model—a classical epidemic framework describing a spreading process where each node can be in one of three states: susceptible (S), infected (I), or recovered (R). At each time step, an infected node infects its susceptible neighbors with probability $\gamma$ and recovers with probability $\delta$, after which it cannot be infected again. The process continues until no infected nodes remain. This model captures the key dynamics of contagion and recovery, which are also relevant to information diffusion and network resilience. Therefore, it provides a reliable basis for evaluating the spreading ability of nodes and for training deep models to understand real-world network dynamics. In our experiments, each node $v$ is initialized as the only infected node, while all others are susceptible. The final number of infected nodes when the process stabilizes is denoted as $F_{v}$, representing the spreading influence of node $v$. The average value of $F_{v}$ over 100 simulations is used as the influence score, denoted by $\bm{Y}_{v}$. Then we run the SIR model for each node in the network to obtain the influence score variable $\bm{Y}$.
@@
where the adjacent matrix $\bm{W}$ is required to satisfy the constraint in Eq. (18). $g_{1}$ and $g_{2}$ can be regarded as two nonlinear functions used to describe the encoding and decoding process, respectively. Therefore, it is noted here that the causal node embedding module can be seen as a graph autoencoder.
@@
4.3 Causal ranking prediction module
@@
Task-aware ranking loss function. We introduce a new ranking loss function, CausalListMLE, which can be viewed as an extension of ListMLE [xia2008listwise] for optimizing the overall ranking prediction module. This list-wise ranking loss directly maximizes the likelihood of the entire ranked permutation, rather than focusing on individual scores as in point-wise losses like MSE. By modeling the ranking distribution, CausalListMLE focuses on preserving the correct relative order among all items, and typically achieves better performance in ranking-based learning tasks. The causal ranking loss $\mathcal{L}_{2}$ is defined as
@@ 4.4 The objective function
To ensure that the learned $\bm{W}$ forms a DAG, this objective must satisfy the acyclicity constraint [zheng2018dags]:
@@
b) (Necessity): Conversely, if $\mathbb{G}$ is acyclic, Lemma 1 implies $q_{ii}^{(k)}=0$ for all $i$ and $k$. Thus, $\mathrm{tr}(\bm{Q}^{k})=0$, and therefore
@@
To solve above equality-constrained problem, the augmented Lagrangian method [ng2019graph] is employed to convert it into an unconstrained problem. After introducing the Lagrange multiplier and the penalty term, the objective function Eq. (4.4) is reformulated as the following augmented Lagrangian function:
@@
4.5 Computational complexity
The computational complexity of ICAN primarily stems from its causal node embedding module and causal ranking prediction module. For the autoencoder in the causal node embedding module, the complexity of encoder is $O(|E|d+nd^{2})$ and the complexity of decoder is $O(n^{2}d)$, resulting in an overall complexity for the Autoencoder of $O(|E|d+nd^{2}+n^{2}d)$, where $n$ and $d$ denote the number of nodes and feature dimensions of the input $\bm{X}$. $\lvert E\rvert$ is the number of edges. As for the causal ranking prediction module, the complexity of the GNN layer is $O(|E|d+nd^{2})$ and the complexity of the FC layer is $O(nd)$. Therefore, the total complexity of the causal ranking prediction module is $O(|E|d+nd^{2})$. Combining both components, the overall computational complexity of the proposed model is $O(|E|d+nd^{2}+n^{2}d)$.
5 Experiment
In this section, experiments are performed on the real-world networks to validate the effectiveness of the proposed ICAN method in the node importance ranking problem. In the following, the experimental settings such as the used datasets, comparison methods, evaluation criterion and model parameters for the experiments are described first, followed by a detailed analysis of the results. To systematically evaluate ICAN, we define three research questions (RQs)—”Three CANs Test for ICAN”:
  • RQ1: Can ICAN trained without reliance on the target network learn causally relevant embeddings of node importance and generalize effectively across diverse real-world network structures?
  • RQ2: Can the causal node embedding mechanism, the feature–task co-optimization mechanism, and the causal ranking loss function in ICAN improve the cross-network ranking performance?
  • RQ3: Can ICAN achieve comparable performance when trained on different types of generative networks, indicating robustness to the network generation process?
5.1 Description of datasets
@@
Training datasets. Five representative synthetic network models are used for training, including Barabási–Albert (BA) [barabasi2009scale], Extreme Homogeneous (EH) [lou2020towards], Erdös-Rényi (ER) random-graph [erdHos1961strength], Q-Snapback (QS) [lou2018toward] and Random Hexagon (RH) [lou2018toward]. The BA networks are generated using the preferential attachment mechanism, where new nodes are sequentially added and connected to existing nodes with a probability proportional to their current degrees, leading to a scale-free degree distribution. The EH networks are obtained by performing additional random edge rectifications on ER networks to transform their degree distribution from Poisson to near-uniform. The ER random graphs are generated by connecting each pair of nodes independently with a fixed probability $p$, producing networks with a binomial degree distribution. The QS networks consist of a directed backbone chain with multiple probabilistic “snapback” edges linking newly added nodes to previously added ones within a defined range, thereby enhancing backward connectivity and robustness. The RH networks are composed of randomly connected hexagonal substructures that capture local clustering and spatial organization similar to lattice-like systems. All trained networks comprise 1,000 nodes with an average degree of 4.
@@
  • Karate [zachary1977information]: A human social network, in which each node represents a member in the club and each link shows the relationship between two members.
  • Jazz [gleiser2003community]: A human collaboration network, in which each node represents a jazz musician and each link indicates that two musicians have performed together in the same band.
  • Email-univ [guimera2003self]: An email communication network from the University Rovira i Virgili in Spain, where nodes represent users and edges indicate at least one email was exchanged.
  • USAir [kunegis2013konect]: A directed weighted network represents the flight connections among 1,574 U.S. airports in 2010. Each node corresponds to an airport, and each directed edge indicates the existence of at least one flight from the origin to the destination airport. The edge weight quantifies the total number of flights operated on that route during the year.
  • Vidal [rual2005towards]: A network represents an initial version of a proteomescale map of Human binary protein–protein interactions.
  • Email-dnc [rossi2015network]: A directed unweighted network derived from the 2016 Democratic National Committee (DNC) email leak. Each node represents an individual user, and each directed edge indicates that at least one email was sent from one user to another.
In our experimental setup, the model is trained exclusively on synthetic networks and evaluated on real-world networks. This setting requires the model to generalize across graphs with different numbers of nodes. To achieve this, we keep the second dimension of the feature matrix $\bm{X}$ (i.e., the node feature dimension) the same for both the training and test graphs. In this way, the input dimension of the GCN weight matrix $\bm{w}$ in the node-wise update function $\mathrm{ReLU}(\widetilde{\bm{A}}\bm{X}\bm{w}+\bm{b})$ remains unchanged between training and testing. The fixed dimensionality of $\bm{w}$, together with parameter sharing across all nodes, allows the same set of parameters to be directly applied to graphs of different sizes. These design choices endow our model with strong inductive generalization capabilities, enabling robust performance in cross-network evaluation scenarios.
5.2 Comparative methods
To demonstrate the advantages of the ICAN method in node ranking, we compare it against a range of state-of-the-art baselines widely used for measuring node importance in complex networks. Specifically, the comparison includes five traditional topology-based methods—encompassing neighborhood-based, eigenvector-based, and path-based techniques—along with six recently developed deep learning-based approaches. A brief summary of these methods is provided below.
  • Degree Centrality (DC) [bonacich1972factoring]. It quantifies the influence of a node based on the number of direct connections it has to other nodes in the graph. The degree centrality of a node u is formally defined as
@@ 5.2 Comparative methods
  • Betweenness Centrality (BC) [freeman1977set]. It measures the importance of a node by quantifying how frequently it appears on the shortest paths between all pairs of nodes in the network.
@@
  • Eigenvector Centrality (EC) [bonacich1987power]. It assigns each node a centrality score based not only on its own connectivity but also on the importance of the nodes it is connected to. The EC score $\bm{X}_{u}$ of node $u$ satisfies
@@
  • $H$-index (HI) [hirsch2005index]. The $H$-index of a node considers the degrees of its neighbors and reflects how many of them are themselves influential. For a node $u$ with neighbors $j_{1},j_{2},\ldots,j_{k}$, the $H$-index is defined as
@@
  • $K$-Shell (KS) [kitsak2010identification]. The $K$-shell algorithm operates through an iterative pruning process. It iteratively peels away network layers to assign each node a $k$-value indicating its coreness based on residual connections. A higher $K$-shell value indicates greater node influence.
  • GNN-Bet [maurya2021graph]. This method enables node ranking by employing a dual-path aggregation mechanism that separately propagates features along incoming and outgoing shortest paths using row-wise modified adjacency matrices, with final ranking scores generated through multiplicative combination of the aggregated path representations.
  • GNN-Close [maurya2021graph]. This approach achieves node ranking through a hierarchical feature aggregation scheme that combines normal adjacency operations in the initial layer with column-wise modified matrices in subsequent layers to constrain feature propagation along shortest paths, producing final ranking scores via summation of layer-wise outputs.
  • RCNN [yu2020identifying]. This method leverages convolutional neural networks to extract features from node-centric subgraph structures and ranks node importance based on learned representations.
  • CGNN [zhang2022new]. This method enhances RCNN by incorporating additional GNN layers, improving the capture of topological dependencies for more accurate vital node identification.
  • AGNN [xiong2024vital]. This method combines a GCN-based autoencoder with a GNN ranking head to jointly learn structural embeddings and predict node importance using listwise ranking loss.
  • STGCN [huang2025sparse]  [^1] Since AGNN and STGCN employ the adjacency matrix as the input to their autoencoder and do not explicitly address the issue of inconsistent numbers of nodes between training and testing networks, we instead incorporate the generated node features by the adjacency matrix as the input to the AGNN and STGCN models.. This method leverages a hybrid architecture combining Sparse Transformer layers for global dependency modeling and GCNs for local topological features to rank key nodes in complex networks.
5.3 Evaluation criterion
In this section, we introduce the evaluation metric used to measure the effectiveness of the predicted ranking results by ICAN, Kendall’s $\tau$ coefficient [knight1966computer]. This non-parametric statistic is widely employed to measure the ordinal association between two ranking lists. It quantifies the similarity between two ranking orders by comparing the numbers of concordant and discordant pairs. The definition of $\tau$ is
$\tau=\frac{2(N_{c}-N_{d})}{n(n-1)}$},
5.4 Implementation details